# Large-scale, Independent and Comprehensive study of the power of LLMs for test case generation

### v1

@@ 5. Threats to Validity
External Validity. The generalizability of our findings to real-world software testing scenarios might be limited by several factors. Our chosen benchmarks (Defects4J, SF110, and a custom dataset) may not fully capture the diverse complexities encountered in real-world software. Additionally, overestimation of LLM capabilities could occur if the models were trained on test cases included within the datasets. We mitigated this risk by creating a novel dataset, but further investigation is warranted. Focusing on four instruct-tuned LLMs from two families (OpenAI and MistralAI) restricts the representativeness of our results to the broader landscape of available LLMs. Furthermore, limitations exist regarding knowledge cutoff dates for the Mistral models. While the official paper lacks details, external sources suggest a 2021-2022 cutoff for Mistral 7B (Otterly, 2024; Saket3199, 2024). For Mixtral 8x7B, we only have the release date (December 2023), which may not reflect its actual knowledge cutoff. It’s crucial to acknowledge that LLMs are constantly evolving. Our findings represent their capabilities only during the study period (November 2023 to April 2024).
Internal Validity. Threats to internal validity in our study stem from the randomness of LLMs and EvoSuite, prompt design, and our bug detection methodology. To address randomness, we conducted each experiment thirty times and employed detailed statistical analysis to ensure reliable interpretations. Our custom-designed prompts, while based on model guidelines and previous research, might not be perfectly optimized for each model or test generation task. Our bug detection methodology mainly focuses on functional correctness, which may miss performance or security-related bugs. We also need to consider flaky tests, as they can result in misleading bug detection outcomes.
@@ 6. Related work
Siddiq et al. (Siddiq et al., 2023) evaluated the ability of LLMs like Codex, GPT-3.5-Turbo, and StarCoder to generate JUnit tests for Java, using benchmarks from HumanEval and EvoSuite SF110. They used zero-shot prompts and focused on aspects like compilation rates, correctness, coverage, and test smells. Our research aligns with their findings, highlighting similar challenges faced by LLMs. However, we further explored prompt engineering techniques to enhance test quality and coverage. Additionally, our study investigated bug detection capabilities, offering a more comprehensive view of LLMs’ applications in software testing.
Tang et al. (Tang et al., 2024) recently evaluated the effectiveness of ChatGPT and EvoSuite in generating Java unit test suites, focusing on correctness, readability, code coverage, and bug detection. Similarly to us, they find ChatGPT underperforms compared to EvoSuite in test suite effectiveness, especially in terms of code coverage. Surprisingly, they have claimed that EvoSuite test suites finds more bugs than LLM-generated test suites. This finding contradicts with our study (on slightly different projects’ code), potentially due to the methodology for bug detection: several works in the literature conclude on bug detection as soon as a test execution fails. Our definition is more constrained. Our study further goes beyond ChatGPT and EvoSuite, assesses various prompt designs, and evaluates on a comprehensive set of testing metrics.
Taking into account the limitations of existing LLMs, Rao et al. (Rao et al., 2023) have proposed to train a specific GPT-style language model, CAT-LM, for the task of test generation. To that end they align code and test files during training. With 2.7 billion parameters, CAT-LM excels in generating syntactically valid tests that closely mimic developer-written tests in structure and style. Unfortunately, CAT-LM has not explored the potential of prompt design in improving the capabilities of LLMs and has not targeted bug detection as a performance criterion.
@@ 7. Conclusion
We conducted a large-scale study investigating LLMs for unit test generation. Our experimental results record for the first time in the literature on applying LLMs for test generation to what extent prompt design significantly impacts test quality: chain-of-thought prompting techniques and our newly designed Guided Tree-of-though prompting strategy yield the most effective tests. While correctness issues persist, compilable LLM-generated tests resembled human-written ones in terms of adherence to coding styles, suggesting potential. A trade-off emerged: EvoSuite achieved higher coverage, while CoT-prompted GPT-3.5 Turbo and Mistral 7B excelled in bug detection. However, this success was primarily observed on known datasets, hinting at some limitations of LLMs with unfamiliar code. These promising yet preliminary results call for further research by the software engineering community to unlock LLMs’ full potential, particularly in generating tests that are relevant to the ultimate task of bug detection.
Open Science To promote transparency and facilitate reproducibility, we make our artifacts available to the community at:

### v2

@@ 5. Threats to Validity
External Validity. The generalizability of our findings to real-world software testing may be limited by several factors. Our benchmarks (Defects4J, SF110, and CMD) may not fully capture the complexities of real-world software. The CMD dataset, with only two projects, limits diversity and representativeness. This small sample size may affect results, particularly for code coverage, where GPT-3.5 Turbo performed better than GPT-4 on CMD—a result that requires cautious interpretation. Further analysis across larger datasets is needed to confirm trends. Additionally, overestimation of LLM capabilities could occur if models were trained on test cases within the datasets, though we mitigated this with CMD. Focusing on four instruct-tuned LLMs from two families (OpenAI and MistralAI) also limits representativeness. Furthermore, the knowledge cutoff dates for Mistral models may affect performance, as Mistral 7B has a 2021-2022 cutoff, and Mixtral 8x7B was released in December 2023. Our findings reflect the models’ capabilities during the study period (November 2023 to April 2024).
Internal Validity. Internal threats stem from the randomness inherent in LLM outputs and EvoSuite, prompt design, and the test smell detection methodology. To account for randomness, we repeated each experiment 30 times and applied statistical analysis to ensure reliable interpretations. While our prompts were designed based on model guidelines and prior research, they may not be fully optimized for every model or task. Our test smell detection methodology, although comprehensive, may miss other quality issues such as performance or security concerns within the test suites. We also need to consider flaky tests, which can result in inconsistent test smell detection outcomes.
@@ 6. Related work
Siddiq et al. (Siddiq et al., 2023) evaluated LLMs like Codex, GPT-3.5-Turbo, and StarCoder for generating JUnit tests, focusing on compilation rates, correctness, coverage, and test smells. Our study aligns with their findings on challenges like persistent test smells, including Magic Number Test and Assertion Roulette. However, we further explored the impact of various prompt techniques to improve test quality and reduce smells. Additionally, we assessed a broader range of LLMs and prompt strategies, providing a more comprehensive analysis of test quality.
Tang et al. (Tang et al., 2024) evaluated ChatGPT and EvoSuite for generating Java unit tests, focusing on correctness, readability, and coverage. Like our study, they found that ChatGPT underperforms in coverage compared to EvoSuite, though it showed potential in generating diverse test inputs. EvoSuite consistently achieved higher coverage overall. Our study extends this by assessing multiple prompt techniques and focusing on test smell detection, offering a deeper understanding of the maintainability and quality of LLM-generated tests.
Alshahwan et al. (Alshahwan et al., 2024) introduced TestGen-LLM, a tool at Meta that improves existing human-written tests using LLMs. Unlike our study, which focuses on generating tests from scratch with various prompts, TestGen-LLM enhances tests by increasing coverage without introducing regressions. During Meta’s test-a-thons, the tool had a 73% acceptance rate for production deployment. While we focus on prompt engineering and test smell detection, both approaches aim to generate high-quality tests using LLMs.
7. Discussion of Findings and Implications
Our findings indicate several key implications for both researchers and developers. The prompt technique plays a crucial role in the quality of LLM-generated test suites. CoT and GToT generally produce better results, indicating that refined prompt designs lead to improved outputs. While LLMs like GPT-3.5 and GPT-4 show potential, especially with CoT, they still lag behind EvoSuite in terms of coverage. However, LLMs excel in generating more readable tests when using well-crafted prompts, suggesting they could complement tools like EvoSuite in scenarios prioritizing readability. A key issue is the persistence of test smells like Magic Number Test across all models and techniques, highlighting the need for human intervention to improve test maintainability. GPT-3.5 Turbo’s better performance over GPT-4 on CMD suggests differences in handling novel codebases, though CMD’s small size calls for cautious interpretation, with larger datasets needed for validation.
@@ 8. Conclusion
We conducted a large-scale study on the effectiveness of LLMs for unit test generation. Our experimental results show, for the first time in the literature, the substantial impact of prompt design on the quality of LLM-generated unit tests. Techniques such as Chain-of-Thought (CoT) and our newly introduced Guided Tree-of-Thought (GToT) prompting strategy produced the highest-quality tests. Although correctness issues persist, compilable LLM-generated unit tests closely resembled human-written ones in terms of adherence to coding standards, highlighting their potential. A key trade-off emerged: while EvoSuite consistently achieved higher code coverage, CoT-prompted GPT-3.5 Turbo and Mistral 7B excelled in generating more readable and maintainable tests. However, these advantages were primarily observed on familiar datasets, pointing to potential limitations in handling unfamiliar codebases. These preliminary, yet promising, results suggest that further research is needed to fully unlock the potential of LLMs in unit test generation, particularly in improving test quality by addressing persistent issues like test smells.
9. Data Availability
To promote transparency and facilitate reproducibility, we make our artifacts available to the community at: